# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ 7.2.2. Methods Targeting the Graph-centric Tasks
RAGRAPH (ref:ragraph, ) performs node classification, link prediction, and graph classification tasks. By retrieving and leveraging the most relevant subgraphs, RAGRAPH passes information within subgraphs to enhance the representation of the center node, facilitating various graph learning tasks. LitFM (ref:litfm, ) considers link prediction (citation link prediction, paper recommendation) and text generation (title generation, abstract completion, citation sentence generation) tasks.
LLMs-EP (ref:Graph-LLM, ) focuses on understanding how text-attribute GNNs can leverage LLMs to enhance node classification tasks. This paper presents an empirical study of two distinct pipelines for incorporating LLMs into graph tasks. In the first pipeline, LLMs are employed as enhancers targeting the text content associated with each node. Specifically, LLMs-EP utilizes LLMs either as embedding encoders to generate node representations or to incorporate additional textual content from related nodes. These approaches are termed feature-level and text-level enhancement, respectively, enabling more informative representations of nodes. In the second pipeline, LLMs-EP provides an in-depth analysis of two scenarios: first, exploring whether LLMs can directly predict node categories, and second, examining the impact of using 2-hop neighbors to incorporate structural information into the prediction process. By combining text and structural data, LLMs-EP aims to improve the accuracy and understanding of node classification in graph-based tasks.
GraphGPT (ref:GraphGPT, ) aims to align LLMs with graph learning tasks through a graph instruction tuning paradigm. The framework focuses on integrating graph structural knowledge with LLMs to enhance generalization in both zero-shot and supervised settings. GraphGPT includes a text-graph grounding component that aligns graph structures with natural language, allowing LLMs to comprehend complex graph structures. This alignment is achieved through a dual-stage instruction tuning process: first, self-supervised instruction tuning is employed to provide structural knowledge, and second, task-specific instruction tuning is used to enhance adaptability across various graph tasks, including node classification and link prediction. GraphGPT also leverages COT distillation to improve step-by-step reasoning abilities, making it effective at addressing the challenges of zero-shot learning in graph-based tasks.
@@ 7.3. Domain-specific Tasks
Healthcare. Existing literature leverages RAG to efficiently summarize long-form medical documents, enhancing LLMs’ ability to generate accurate and evidence-based responses. For instance, MedGraphRAG (ref:medgraphrag, ) serves as a dedicated framework for utilizing graph-based RAG in healthcare applications. Its retrieval process involves generating a tag-summary on the user query, identifying the most relevant graph through similarity-based selection, and refining responses. HyKGE (ref:hykge, ) is a framework leveraging RAG based on knowledge graphs in LLM-empowered medical applications. To improve medical consultation quality, HyKGE integrates a granularity-aware reranking module to eliminate noise while preserving diversity-relevance balance in retrieved knowledge. REALM (zhu2024realm, ) enhances the clinical predictive capabilities by integrating RAG with multimodal Electronic Health Records (EHR). It enhances the utilization of EHR data in healthcare and reconciles it with the intricate medical context necessary for educated clinical forecasts.
@@
Scientific Literature. By harnessing the capabilities of GraphRAG, it becomes possible to navigate the vast expanse of scientific literature, extracting key insights and condensing complex ideas and high-quality knowledge. For example, DALK (ref:dalk, ) constructs a specific knowledge graph sourced from Alzheimer’s Disease scientific literature. The essential knowledge is filtered through a coarse-to-fine sampling algorithm so that it can offer valuable insights. ATLANTIC (ref:atlantic, ) introduces a retrieval-augmented language model that incorporates structural awareness for interdisciplinary scientific tasks, integrating heterogeneous document graphs to capture structural relationships among scientific documents. By fusing textual and structural embeddings, ATLANTIC enhances the retrieval of coherent and faithful passages. KGP (ref:kgp, ) creates a KG over multiple kinds of literature and employs an LLM-guided traversal agent to retrieve and synthesize relevant information for answering complex queries.
@@
Code Completion. In code completion tasks, using a graph with the RAG system helps capture and represent the complex relationships between code elements, such as control flow, data dependencies, and function calls. For example, GraphCoder (ref:graphcoder, ) is an RAG framework designed to improve code completion by integrating repository-specific knowledge into code LLMs. Unlike traditional methods, it uses a Code Context Graph (CCG), which captures relationships like control flow and data dependence. The framework employs a two-step retrieval process: coarse-grained filtering to find candidate code snippets, followed by fine-grained re-ranking to prioritize those with aligned dependencies. This combination of structural and lexical context leads to more relevant code snippets, improving the accuracy of code generation. GraphCoder outperforms baseline methods, achieving higher code and identifier exact matches with lower time and storage costs. Its language-agnostic design, tested on Python and Java, proves its versatility across different programming environments.
@@
Biomedical Question Answering. Biomedical knowledge is specialized and constantly changing, making it hard for traditional language models to provide accurate, up-to-date answers without access to external, structured information. GLBK (ref:noname5, ) aims to tackle the information overload problem in biomedical question answering. The primary objective is to retrieve and prioritize relevant biomedical documents from a vast corpus, such as PubMed, while addressing complex, open-ended queries like identifying drug targets for diseases. Unlike traditional methods relying solely on embedding similarity, this approach integrates knowledge graph structures to mitigate biases from overrepresented topics and improve access to the long tail of biomedical knowledge. By leveraging entity recognition, relationship extraction, and graph-based rebalancing, the method enhances retrieval precision and ensures more contextually relevant information is surfaced.
@@ 8. Future Research Opportunities
Adaptive Prompts for LLMs. Future research could delve into the design and evaluation of adaptive prompts to better align with the pretrained knowledge of LLMs. By tailoring the structure, terminology, and hierarchy of information drawn from graphs, system designers could harness improved synergy between graph-based data and the LLM’s reasoning capabilities. One could also investigate methods to prioritize the most critical entities, relationships, or paths in a query (ref:gnn-ret, ; ref:kg-rank, ), ensuring that the prompt remains precise and contextually rich without overwhelming the model. Such selective attention could enhance both the relevance and interpretability of LLM outputs. Furthermore, exploring feedback mechanisms that dynamically refine the prompt based on the LLM’s responses might be an interesting direction, allowing systems to continuously optimize for coherence and clarity. Ultimately, adaptive prompt strategies could pave the way for more accurate, context-aware question answering, harnessing the potential of complex graph structures in tandem with cutting-edge language models.
@@
Enhanced Understanding for Graph Problems. Future research could delve into approaches that enable deeper structural comprehension of graphs within LLMs. Since many current models rely on sequential token representations, they struggle to handle larger or more complex graph structures—especially when a sequence involves a large number of nodes (graphwiz, ; ref:GPT4Graph, ). This limitation hinders their ability to solve typical graph-centric problems, such as path finding or community detection, which often require more specialized, non-linear reasoning. One direction might be to integrate graph-specific modules or adopt novel representations that faithfully capture relational information without overwhelming the model. Moreover, another avenue of exploration could involve combining LLMs with algorithmic components adopted at handling extensive, node-rich graphs.
Dynamic Graphs for RAG. A promising future direction for RAG lies in integrating dynamic and temporal graphs, addressing the limitations of static graph-based retrieval systems. Currently, RAG systems of graphs mainly rely on static structure for knowledge retrieval, but in real-world applications, graphs are inherently dynamic, evolving over time as new information becomes available (IDOL, ; quota, ). Recent studies, such as DynaGRAG (DynaGRAG, ), RAG4DyG (RAG4DyG, ), DRAGIN (DRAGIN, ), and Multi-Armed Bandit Enhanced RAG (MAB-RAG, ), have begun exploring dynamic retrieval mechanisms and temporal-aware modeling to address these limitations. However, these approaches still exhibit certain drawbacks, including significant computational overhead, dependency on the quality of dynamically retrieved data, and scalability challenges. Their effectiveness in capturing real-world temporal dynamics remains limited, highlighting the need for further advancements. One major challenge of developing efficient RAG systems on dynamic graphs is that the retrieval process needs to be repeated frequently to capture the most recent updates once the updates enter the system, leading to significant computational overhead (agenda, ). Additionally, current RAG systems fail to incorporate the temporal information of the graph, which is crucial in fields like healthcare and economics. For instance, in healthcare, the relationships between diseases, treatments, and outcomes evolve over time, and static graphs fail to reflect the latest medical research or patient data (gao2022modeling, ). Similarly, in economics, market conditions, supply chains, and consumer behavior are constantly changing (harrod1973economic, ), and failing to account for these temporal changes can result in outdated or inaccurate information retrieval. Addressing these challenges by integrating temporal graphs into RAG systems will enhance their scalability, efficiency, and relevance, enabling more accurate and timely predictions, particularly in rapidly changing domains.
@@
Multi-Modal Graphs for Cross-Domain RAG. Future advancements in Multi-Modal Graphs for Cross-Domain RAG should focus on the development of adaptable retrieval algorithms that accommodate the diversity of data types encountered in real-world applications. Presently, graph-based RAG systems are largely confined to single-domain use cases, often employing retrieval methods tailored to specific types of textual information. However, as cross-domain scenarios become increasingly common, there is a pressing need for retrieval techniques that can seamlessly handle multi-modal graph structures, including images (han2022vision, ), audio (nie2020c, ), and numerical data. These enhanced retrieval algorithms must not only integrate and interpret a wide range of data formats but also dynamically adjust their retrieval strategies to match the unique characteristics of each domain. By enabling a more flexible, multi-modal approach, future research in this field can overcome the limitations of current systems, paving the way for more robust and versatile knowledge acquisition across diverse domains.
Scalable Graphs Retrieval in RAG. As knowledge graphs grow more expansive and intricate, traditional RAG methods struggle to maintain performance, especially with respect to processing speed and resource usage efficiency. To overcome this, future research may focus on developing advanced indexing techniques and graph partitioning strategies that enable faster query processing across distributed systems such as the methods employed in (liu2022net, ). Additionally, the graph retrieval methods leveraging GNNs for efficient representation learning should also consider the scalability of graph embeddings (liao2022scara, ; liu2023simga, ), enabling more accurate and faster retrieval even from large-scale graphs. Another key area will be optimizing algorithms for real-time graph updates, ensuring that retrieval processes remain effective as new information is continuously integrated into the graph.
@@
Improved Graph Construction Techniques. The future of RAG systems can center around overcoming the limitations of traditional triple-based knowledge graphs, which are often too simplistic for complex, real-world data  (ref:noname8, ). While the subject-predicate-object (zhang2017visual, ) format has served as a foundation, it does not fully capture intricate relationships and evolving patterns within knowledge. To address this, future graph construction methods should explore hypergraphs, which allow for multi-node relationships (li2007network, ), and semantic embeddings (wang2018zero, ) that can represent context, enhancing the depth and accuracy of graph structures. Additionally, GNNs can be used to learn graph representations that dynamically evolve as new information is introduced, allowing for more precise and adaptive knowledge storage. Another important direction will be the development of hierarchical graphs such as the approaches utilized in (ying2018hierarchical, ) that capture different levels of abstraction, providing a more flexible and structured way to organize knowledge. Furthermore, adaptive graph construction techniques, which adjust the structure and organization of the graph based on the type and volume of data, will be essential for handling real-time updates without sacrificing retrieval speed. In summary, these advanced techniques will allow RAG systems to store knowledge more efficiently, retrieve information faster, and support more complex, multi-modal data interactions.
Explainability in Graph-Based RAG. The retrieval and prompting process in RAG systems currently operates as a black box, which can undermine trust and transparency. Users have little visibility into how knowledge is retrieved or why certain information is selected, leading to doubts about the reliability of the system. To improve this, explainability techniques must be integrated into graph-based RAG systems to improve the interpretability of the decision-making process. One key challenge is explaining how the graph’s structure and the relationships between nodes influence the retrieval results. For example, providing visualizations of the graph’s relevant nodes and edges could help users understand which parts of the knowledge graph contributed to a particular retrieval. Additionally, interpretability frameworks that offer clear, step-by-step breakdowns of how the graph responds to user queries would build greater trust in the system. By improving the transparency of these processes, users can gain more confidence in the decisions made by RAG systems, making them more reliable and accessible, especially in high-stakes applications.

### v2

@@ 7.2.2. Methods Targeting the Graph-centric Tasks
RAGRAPH (Jiang et al., 2024b) performs node classification, link prediction, and graph classification tasks. By retrieving and leveraging the most relevant subgraphs, RAGRAPH passes information within subgraphs to enhance the representation of the center node, facilitating various graph learning tasks. LitFM (Zhang et al., 2024a) considers link prediction (citation link prediction, paper recommendation) and text generation (title generation, abstract completion, citation sentence generation) tasks.
GraphGPT (Tang et al., 2024) aims to align LLMs with graph learning tasks through a graph instruction tuning paradigm. The framework focuses on integrating graph structural knowledge with LLMs to enhance generalization in both zero-shot and supervised settings. GraphGPT includes a text-graph grounding component that aligns graph structures with natural language, allowing LLMs to comprehend complex graph structures. This alignment is achieved through a dual-stage instruction tuning process: first, self-supervised instruction tuning is employed to provide structural knowledge, and second, task-specific instruction tuning is used to enhance adaptability across various graph tasks, including node classification and link prediction. GraphGPT also leverages COT distillation to improve step-by-step reasoning abilities, making it effective at addressing the challenges of zero-shot learning in graph-based tasks.
@@ 7.3. Domain-specific Tasks
Healthcare. Existing literature leverages RAG to efficiently summarize long-form medical documents, enhancing LLMs’ ability to generate accurate and evidence-based responses. For instance, MedGraphRAG (Wu et al., 2024b) serves as a dedicated framework for utilizing graph-based RAG in healthcare applications. Its retrieval process involves generating a tag-summary on the user query, identifying the most relevant graph through similarity-based selection, and refining responses. HyKGE (Jiang et al., 2024c) is a framework leveraging RAG based on knowledge graphs in LLM-empowered medical applications. To improve medical consultation quality, HyKGE integrates a granularity-aware reranking module to eliminate noise while preserving diversity-relevance balance in retrieved knowledge. REALM (Zhu et al., 2024b) enhances the clinical predictive capabilities by integrating RAG with multimodal Electronic Health Records (EHR). It enhances the utilization of EHR data in healthcare and reconciles it with the intricate medical context necessary for educated clinical forecasts.
@@
Scientific Literature. By harnessing the capabilities of GraphRAG, it becomes possible to navigate the vast expanse of scientific literature, extracting key insights and condensing complex ideas and high-quality knowledge. For example, DALK (Li et al., 2024a) constructs a specific knowledge graph sourced from Alzheimer’s Disease scientific literature. The essential knowledge is filtered through a coarse-to-fine sampling algorithm so that it can offer valuable insights. ATLANTIC (Munikoti et al., 2023) introduces a retrieval-augmented language model that incorporates structural awareness for interdisciplinary scientific tasks, integrating heterogeneous document graphs to capture structural relationships among scientific documents. By fusing textual and structural embeddings, ATLANTIC enhances the retrieval of coherent and faithful passages. KGP (Wang et al., 2024d) creates a KG over multiple kinds of literature and employs an LLM-guided traversal agent to retrieve and synthesize relevant information for answering complex queries.
@@
Code Completion. In code completion tasks, using a graph with the RAG system helps capture and represent the complex relationships between code elements, such as control flow, data dependencies, and function calls. For example, GraphCoder (Liu et al., 2024e) is an RAG framework designed to improve code completion by integrating repository-specific knowledge into code LLMs. Unlike traditional methods, it uses a Code Context Graph (CCG), which captures relationships like control flow and data dependence. The framework employs a two-step retrieval process: coarse-grained filtering to find candidate code snippets, followed by fine-grained re-ranking to prioritize those with aligned dependencies. This combination of structural and lexical context leads to more relevant code snippets, improving the accuracy of code generation. GraphCoder outperforms baseline methods, achieving higher code and identifier exact matches with lower time and storage costs. Its language-agnostic design, tested on Python and Java, proves its versatility across different programming environments.
@@
Biomedical Question Answering. Biomedical knowledge is specialized and constantly changing, making it hard for traditional language models to provide accurate, up-to-date answers without access to external, structured information. GLBK (Delile et al., 2024) aims to tackle the information overload problem in biomedical question answering. The primary objective is to retrieve and prioritize relevant biomedical documents from a vast corpus, such as PubMed, while addressing complex, open-ended queries like identifying drug targets for diseases. Unlike traditional methods relying solely on embedding similarity, this approach integrates knowledge graph structures to mitigate biases from overrepresented topics and improve access to the long tail of biomedical knowledge. By leveraging entity recognition, relationship extraction, and graph-based rebalancing, the method enhances retrieval precision and ensures more contextually relevant information is surfaced.
7.4. Evaluation of Tasks
Open-source benchmark datasets. Following the taxonomy above, we also group the open-source benchmark datasets into three representative families (summarized in Table 6). For KGQA, WebQ (Berant et al., 2013) and WebQSP (Yih et al., 2016) focus on entity linking and compositional queries over knowledge graphs, while HotpotQA (Yang et al., 2018) emphasizes multi-hop aggregation and GrailQA (Gu et al., 2021) targets schema generalization across unseen relations. Recent temporal KGQA corpora—CronQuestion (Saxena et al., 2021), Forecast (Ding et al., 2023), and MultiTQ (Chen et al., 2023)—introduce explicit time constraints to test era-aware retrieval and timestamp-valid answers. For graph-centric tasks, RelBench (Robinson et al., 2024) provides standardized node, edge, and graph prediction suites. TUDataset (Morris et al., 2020) offers diverse graph classification collections, and GLBench (Li et al., 2024d) together with GraphEval2000 (Wu et al., 2024c) concentrate on LLM-oriented graph reasoning and program execution. In domain-specific settings, MedQA (Jin et al., 2020) and PubMedQA (Jin et al., 2019) address clinically grounded QA with domain terminology and long-context evidence. BioRED (Luo et al., 2022) targets biomedical entity and relation extraction, and AlphaFin (Li et al., 2024b) represents financial analysis and forecasting under realistic market dynamics.
Table 6. Evaluation of different tasks.
Tasks Metrics Benchmark datasets
KGQA Tasks Exact Match, F1 Score, SCR, Accuracy, ROUGE, BLEU, METEOR, TTFT SimpleQuestion (Bordes et al., 2015), WebQ (Berant et al., 2013), CWQ (Talmor and Berant, 2018), MetaQA (Zhang et al., 2018), CronQuestion (Saxena et al., 2021), Forecast (Ding et al., 2023), MultiTQ (Chen et al., 2023), GraphQA (He et al., 2024), GRBENCH (Jin et al., 2024b), CRAG (Yang et al., 2024c), NQ (Kwiatkowski et al., 2019), KILT (Petroni et al., 2021), MultiHop (Tang and Yang, 2024), CRUD-RAG (Lyu et al., 2024), CREAK (Onoe et al., 2021), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), FACTKG (Kim et al., 2023b), WebQSP (Yih et al., 2016), GrailQA (Gu et al., 2021),
Graph Tasks Accuracy, F1 Score, MRR, NDCG, Hit Ratio, MAE, RMSE RelBench (Robinson et al., 2024), TabGraphs (Bazhenov et al., 2024), HiTab (Cheng et al., 2021), TUDataset (Morris et al., 2020), GLBench (Li et al., 2024d), GPT4Graph (Guo et al., 2023), CS-TAG (Yan et al., 2023), GraphEval2000 (Wu et al., 2024c)
Domain-specific Tasks Accuracy, F1 Score, MRR, ROUGE, RMSD, Cumulative Return STARK (Wu et al., 2024e), AlphaFin (Li et al., 2024b), MedQA (Jin et al., 2020), BioRED (Luo et al., 2022), QALD (Perevalov et al., 2022), Mintaka (Sen et al., 2022), TutorQA (Yang et al., 2024b), UltraDomain (Qian et al., 2024), PubMedQA (Jin et al., 2019)
Evaluation metrics. We summarize widely used metrics across the three task families and note that Accuracy and F1 score are ubiquitous core measures for them. For answer quality in, Exact Match (EM), token-level F1 score, and Accuracy are standard, with Recall-Oriented Understudy for Gisting Evaluation (ROUGE), Bilingual Evaluation Understudy (BLEU), and Metric for Evaluation of Translation with Explicit ORdering (METEOR) applied when systems produce long-form justifications. For ranking and retrieval in graph reasoning, Mean Reciprocal Rank (MRR), Normalized Discounted Cumulative Gain (NDCG), and Hits@k are additionally reported. In domain-specific tasks, biomedical information extraction typically uses Accuracy and micro or macro F1; biology-oriented structure comparison relies on Root Mean Square Deviation (RMSD); and finance-oriented studies commonly track Cumulative Return, Sharpe Ratio, Maximum Drawdown and so on. Efficiency and stability are often captured with latency measures such as Time to First Token (TTFT) and robustness measures such as Self-Consistency Rate (SCR). The complete introduction of each metric is detailed in the appendix.
7.5. Importance of Graph in Different Task Families
We finally examine the advantages of graph-based RAG over non-graph retrieval across three task families. (a) For simple factoid KGQA, a non-graph retriever often suffices because the answer typically resides in a single passage with little need for disambiguation. For complex KGQA, graphs add clear value by grounding entities and relations, composing multi-hop paths before generation, enforcing type and temporal constraints, constraining search to relevant neighborhoods, and preserving provenance for verification. These capabilities are instantiated by the Prize Collecting Steiner Tree module in G-Retriever, which selects an evidence subgraph aligned to the question, and the Personalized PageRank diffusion module in HippoRAG, which prioritizes neighborhoods seeded by detected entities to guide ranking and multi-hop reasoning. (b) For graph-centric tasks such as node classification and link prediction, performance hinges on local topology, relation types, and neighborhood signals. Graph-based RAG contributes by retrieving a compact subgraph anchored on the target nodes or node pairs, preserving adjacency and type constraints so the generator reasons over dense, topology-aware evidence within a small token budget. This reduces distraction from long passages, improves label and edge consistency, and supports clear provenance. These advantages are exemplified by subgraph message passing modules in RAGRAPH, which strengthen center node representations for node and link prediction, and the feature&text level enhancement modules in LLMs- EP, which inject two-hop structural signals to improve node classification. (c) Domain-specific tasks demand high accuracy and transparent explanations. Graph-based RAG strengthens accuracy by supporting loop and tree pipelines that let the model revisit, branch, and refine retrieval and generation across multiple passes. It also advances explainability by representing structured knowledge beyond classic knowledge graphs, including document, citation, clinical, and code graphs, so the system returns subgraphs with explicit entities, relations, and paths as verifiable evidence. These advantages are illustrated by the traversal agent in KGP, which iteratively expands and consolidates evidence on literature graphs, and the shortest-path retrieval in GLBK, which surfaces indirect biomedical relations with clear provenance.
8. Graph-based RAG in dynamic scenarios
Graph-based RAG in dynamic scenarios increasingly couples time-aware indexing with retrieval that respects event order and temporal constraints. T-GRAG (Li et al., 2025a) constructs a time-stamped knowledge graph and performs layered retrieval that first isolates time-relevant subgraphs, then filters candidate nodes and finally selects fine-grained facts for generation, thereby avoiding cross-time conflicts. DyG-RAG (Sun et al., 2025) defines dynamic event units with explicit temporal anchors and links them into an event graph, retrieving coherent timelines and guiding generation with time-aware chain-of-thought. STAR-RAG (Zhu et al., 2025) compresses temporal regularities into a rule graph and applies seeded personalized PageRank to prioritize a time-aligned neighborhood that matches a query’s temporal intent. TimeRAG (Yang et al., 2025) builds a time series knowledge base and retrieves reference sequences via Dynamic Time Warping before prompting an LLM forecaster with the retrieved patterns. TS-Retriever (Wu et al., 2024a) improves retrieval for time-sensitive questions by training a contrastive, temporally aware retriever with query-side fine-tuning and routing, so that queries with implicit or interval constraints are matched to correctly dated evidence. DRAGIN (Su et al., 2024) triggers retrieval adaptively during generation and formulates queries to reflect the model’s immediate information needs, inserting external knowledge only when beneficial. RAG4DyG (Wu et al., 2024f) retrieves temporally and contextually similar demonstrations from across a dynamic graph and fuses them with the query sequence to adapt to evolving interaction patterns. DynaGRAG (Thakrar, 2024) enhances subgraph representations through de-duplication, two-step pooling, query-aware retrieval, and dynamic similarity–aware traversal to prioritize diverse yet relevant subgraphs as evidence. GenTKG (Gao et al., 2024a) retrieves ordered historical events from a temporal knowledge graph as instruction-style context to support generative forecasting of future links. Collectively, these methods show a convergent design: encode temporal structure explicitly, retrieve evidence that is locally time-consistent, and compose temporally ordered contexts so that generation remains faithful to evolving knowledge.
9. Practical Applications in Industry
Industrial applications of graph-based RAG can be grouped into two tiers: production RAG systems and the workflow–platform stacks. The production tier includes Microsoft’s GraphRAG [^12] https://github.com/microsoft/graphrag, which builds entity-based knowledge graphs and pre-generated community summaries to improve query-focused summarization; NebulaGraph’s GraphRAG [^13] https://www.nebula-graph.io/posts/graph-RAG, which integrates large-language-model reasoning with the NebulaGraph database; Eosphoros’s GraphRAG [^14] https://github.com/eosphoros-ai/DB-GPT, which combines DB-GPT (Zhou et al., 2024a), OpenSPG [^15] https://github.com/OpenSPG/openspg, and TuGraph  [^16] https://tugraph.tech/ for triple extraction, subgraph location, and traversal; Neo4j’s NaLLM [^17] https://github.com/neo4j/NaLLM, which enables natural-language interfaces to knowledge graphs, construction from unstructured data, and report generation; Writer [^18] https://writer.com/, which utilizes graph-based RAG to enhance enterprise content generation and Q&A systems; and Leroy Merlin [^19] https://www.lettria.com/case-study/leroy-merlin-knowledge-graph-product-recommendations, which uses product knowledge graphs to refine recommendation systems, providing personalized customer experiences. The workflow–platform tier comprises LangChain [^20] https://www.langchain.com/ and LangGraph [^21] https://langchain-ai.github.io/langgraph for graph-structured pipelines; LlamaIndex’s Agentic Document Workflows [^22] https://www.llamaindex.ai/blog/agentic-rag-with-llamaindex-2721b8a49ff6 for end-to-end document processing and meta-agent coordination; multi-agent frameworks such as CrewAI [^23] https://github.com/crewAIInc/crewAI and AG2 AutoGen [^24] https://github.com/ag2ai/ag2; and cloud platforms including Vertex AI [^25] https://cloud.google.com/vertex-ai, Amazon Bedrock [^26] https://aws.amazon.com/cn/bedrock/, and IBM watsonx [^27] https://www.ibm.com/products/watsonx.
The Usage of Graph-based RAG in Frontier LLMs. Frontier LLMs such as GPT-5.2  [^28] https://openai.com/index/introducing-gpt-5-2/ and Gemini 3 Pro  [^29] https://ai.google.dev/gemini-api/docs/gemini-3 make graph-based RAG more practical because they are stronger at long-context understanding, tool selection, and multi-step planning. Therefore they can reliably act as controllers and verifiers in retrieval pipelines. In particular, both ecosystems now provide effective retrieval tools named as File Search, which ingest enterprise documents, build searchable indexes, and retrieve only the most relevant evidence for generation  [^30] https://platform.openai.com/docs/guides/migrate-to-responses [^31] https://ai.google.dev/gemini-api/docs/file-search. Under our broad definition, this enables graph-based RAG beyond classic knowledge-graph traversal, where document-level structures and hierarchical graphs can be distilled from text and metadata, and retrieval can return compact subgraphs that enforce entity constraints before generation while reducing token and latency costs. Moreover, Gemini offers explicit controls for extended reasoning, enabling loop and tree style pipelines that iteratively retrieve evidence, verify intermediate results, and revise outputs, which matches the needs of graph-structured pipeline execution in practice  [^32] https://ai.google.dev/gemini-api/docs/thinking. Compared with earlier LLMs, frontier models mainly address two practical bottlenecks in RAG: they reduce controller brittleness for iterative retrieve–verify–revise pipelines, and they make RAG scalable to large private corpora by relying on managed retrieval interfaces instead of manual prompt stuffing (Li et al., 2025b).
Efficiency Discussion. The efficiency of industrial graph-based RAG can be examined through three views. First, graph construction and precomputation determine the baseline latency and storage footprint: Microsoft’s GraphRAG constructs entity-centric knowledge graphs and pre-generates community summaries to speed query-focused summarization at the expense of build time and storage, while Neo4j’s NaLLM automates the extraction of nodes, relations, and properties from unstructured data to reduce manual engineering effort in the indexing stage. Second, the retrieval strategy will control most online computation: AntGroup’s GraphRAG narrows work by locating keywords, mapping them to graph nodes, and traversing only the necessary subgraph with breadth-first or depth-first search, and NebulaGraph’s GraphRAG integrates large language models with the graph database to raise precision and avoid wasteful candidate expansion. Third, pipeline design governs coordination overhead and response quality. Sequential chains centralize control and keep latency low, exemplified by LangChain’s modular RAG chains and CrewAI’s support for sequential processes. Loop pipelines add iterative refinement and self-correction when increasing the cost. For example, LangGraph provides explicit loop capability with state persistence so a retriever and generator can revisit intermediate results when needed. Tree pipelines branch sub-queries or specialized agents for parallel search, which raises coordination cost but can improve answer quality. For instance, AG2 AutoGen together and OpenAI Swarm framework [^33] https://github.com/openai/swarm fans out sub-queries before synthesis, which improves coverage at the cost of additional coordination.
@@ 10. Future Research Opportunities
Adaptive Prompts for LLMs. Future research could delve into the design and evaluation of adaptive prompts to better align with the pretrained knowledge of LLMs. By tailoring the structure, terminology, and hierarchy of information drawn from graphs, system designers could harness improved synergy between graph-based data and the LLM’s reasoning capabilities. One could also investigate methods to prioritize the most critical entities, relationships, or paths in a query (Li et al., 2024f; Yang et al., 2024a), ensuring that the prompt remains precise and contextually rich without overwhelming the model. Such selective attention could enhance both the relevance and interpretability of LLM outputs. Furthermore, exploring feedback mechanisms that dynamically refine the prompt based on the LLM’s responses might be an interesting direction, allowing systems to continuously optimize for coherence and clarity. Ultimately, adaptive prompt strategies could pave the way for more accurate, context-aware question answering, harnessing the potential of complex graph structures in tandem with cutting-edge language models.
@@
Enhanced Understanding for Graph Problems. Future research could delve into approaches that enable deeper structural comprehension of graphs within LLMs. Since many current models rely on sequential token representations, they struggle to handle larger or more complex graph structures—especially when a sequence involves a large number of nodes (Chen et al., 2024a; Guo et al., 2023). This limitation hinders their ability to solve typical graph-centric problems, such as path finding or community detection, which often require more specialized, non-linear reasoning. One direction might be to integrate graph-specific modules or adopt novel representations that faithfully capture relational information without overwhelming the model. Moreover, another avenue of exploration could involve combining LLMs with algorithmic components adopted at handling extensive, node-rich graphs.
@@
Multi-Modal Graphs for Cross-Domain RAG. Future advancements in Multi-Modal Graphs for Cross-Domain RAG should focus on the development of adaptable retrieval algorithms that accommodate the diversity of data types encountered in real-world applications. Presently, graph-based RAG systems are largely confined to single-domain use cases, often employing retrieval methods tailored to specific types of textual information. However, as cross-domain scenarios become increasingly common, there is a pressing need for retrieval techniques that can seamlessly handle multi-modal graph structures, including images (Han et al., 2022), audio (Nie et al., 2020), and numerical data. These enhanced retrieval algorithms must not only integrate and interpret a wide range of data formats but also dynamically adjust their retrieval strategies to match the unique characteristics of each domain. By enabling a more flexible, multi-modal approach, future research in this field can overcome the limitations of current systems, paving the way for more robust and versatile knowledge acquisition across diverse domains.
@@
Improved Graph Construction Techniques. The future of RAG systems can center around overcoming the limitations of traditional triple-based knowledge graphs, which are often too simplistic for complex, real-world data  (Lan and Jiang, 2020). While the subject-predicate-object (Zhang et al., 2017) format has served as a foundation, it does not fully capture intricate relationships and evolving patterns within knowledge. To address this, future graph construction methods should explore hypergraphs, which allow for multi-node relationships (Li and Horvath, 2007), and semantic embeddings (Wang et al., 2018) that can represent context, enhancing the depth and accuracy of graph structures. Additionally, GNNs can be used to learn graph representations that dynamically evolve as new information is introduced, allowing for more precise and adaptive knowledge storage. Another important direction will be the development of hierarchical graphs such as the approaches utilized in (Ying et al., 2018) that capture different levels of abstraction, providing a more flexible and structured way to organize knowledge. Furthermore, adaptive graph construction techniques, which adjust the structure and organization of the graph based on the type and volume of data, will be essential for handling real-time updates without sacrificing retrieval speed. In summary, these advanced techniques will allow RAG systems to store knowledge more efficiently, retrieve information faster, and support more complex, multi-modal data interactions.